# A Chinese Continuous Sign Language Dataset Based on Complex Environments

### v1

@@ Abstract
The current bottleneck in continuous sign language recognition (CSLR) research lies in the fact that most publicly available datasets are limited to laboratory environments or television program recordings, resulting in a single background environment with uniform lighting, which significantly deviates from the diversity and complexity found in real-life scenarios. To address this challenge, we have constructed a new, large-scale dataset for Chinese continuous sign language (CSL) based on complex environments, termed the complex environment - chinese sign language dataset (CE-CSL). This dataset encompasses 5,988 continuous CSL video clips collected from daily life scenes, featuring more than 70 different complex backgrounds to ensure representativeness and generalization capability. To tackle the impact of complex backgrounds on CSLR performance, we propose a time-frequency network (TFNet) model for continuous sign language recognition. This model extracts frame-level features and then utilizes both temporal and spectral information to separately derive sequence features before fusion, aiming to achieve efficient and accurate CSLR. Experimental results demonstrate that our approach achieves significant performance improvements on the CE-CSL, validating its effectiveness under complex background conditions. Additionally, our proposed method has also yielded highly competitive results when applied to three publicly available CSL datasets. Code and dataset is available at https://github.com/woshisad159/TFNet.git.
@@ I Introduction
Sign language, as a unique non-verbal language system, relies on rich hand movements, gesture forms, and body language, serving as the core communication mechanism within the deaf community [1][2][3]. Sign language recognition has always been a focal point for researchers, exemplifying interdisciplinary studies that integrate the essence of computer vision, natural language processing, and human-computer interaction technologies[4][5]. The development of sign language recognition technology has also built a bridge for communication between the deaf and hearing populations. By real-time translation of sign language videos into understandable text or speech, this technology not only simplifies the complexity of cross-language communication but also demonstrates great potential in various fields such as education, employment, and public services[6].
@@
In the field of CSLR, achieving accurate recognition of continuous sign language relies heavily on having a high-quality CSL dataset[7][8][9][10]. Koller et al.[11] constructed the PHOENIX14-2014 dataset by collecting weather forecast programs broadcasted by a German public television station, providing valuable resources for CSLR research. This dataset is based on demonstrations by professional sign language interpreters; however, its collection environment remains limited to the television studio, lacking the diversity of real-world contexts. Camgoz et al.[12] further expanded upon this by introducing the PHOENIX14-2014T dataset, increasing the volume of data and refining the definition of sign boundaries, thereby enhancing annotation accuracy. Although this dataset improves in terms of professionalism, it is still constrained by relatively idealized background conditions, making it difficult to comprehensively reflect the sign language communication scenarios of everyday life. Adaloglou et al.[13], focusing on practical application scenarios, particularly emphasized communication between deaf individuals and public service institutions, such as police departments, hospitals, and citizen service centers. They meticulously designed five typical scenarios, simulating dialogues of varying difficulty levels, thus enhancing the practicality and authenticity of the dataset. However, even with these efforts, the specialized setting of the collection environment and the backgrounds of the participants still fall short of reflecting real life, limiting the representativeness of the dataset and the performance of models in real-world settings. Despite the significant role these datasets have played in advancing CSLR technology, they generally face a common challenge: a disconnect between the collection environment and real-life conditions, as illustrated in Figure 1.
@@
Figure 1: Sample image sequences from four publicly available continuous sign language datasets. These image sequences illustrate that the datasets were collected in laboratory environments or during television programs, with the subjects against a plain-colored background and under uniform lighting conditions.
@@
To overcome the limitations of existing continuous sign language datasets, particularly their disconnection from real-life scenarios, this study constructs a CSL dataset oriented towards practical application environments. The dataset aims to facilitate the seamless transition of CSLR technology from laboratory settings to everyday life, laying a solid foundation for barrier-free communication between the deaf and hearing communities. The construction process of the CE-CSL dataset strictly adheres to the principle of practical application orientation, collecting a large number of continuous sign language video materials derived from real-life scenarios, covering a wide range of situational variations and environmental complexities. After the completion of the CE-CSL dataset, we conducted a comprehensive evaluation using advanced deep learning models to verify its reliability and value as a research benchmark.
Furthermore, this study proposes a TFNet model for continuous sign language recognition. This model, after extracting frame-level features, utilizes information from both the temporal and spectral domains to extract sequential features separately, followed by fusion, ultimately achieving efficient and accurate semantic parsing. Experimental results show that our method achieves significant performance improvements on the CE-CSL dataset, demonstrating its effectiveness under complex background conditions. Additionally, our proposed method has also yielded highly competitive results when applied to three publicly available continuous sign language datasets.
@@
    We have constructed a Chinese Continuous Sign Language dataset CE-CSL dataset oriented towards practical application environments. This dataset, with its rich content and precise annotations, aims to provide a robust foundational resource for the advancement of CSLR technology.
@@
    We propose a TFNet model for CSLR. This model, by leveraging information from both the temporal and spectral domains to extract sequence features, achieves efficient and accurate semantic parsing, significantly improving the accuracy of CSLR in complex environments.
@@
    Our proposed TFNet model demonstrates state-of-the-art performance on the CE-CSL dataset. Additionally, it achieves highly competitive results on three other publicly available CSLR datasets, indicating that TFNet not only excels on the CE-CSL dataset but also exhibits high generality and robustness across different datasets.
@@
Dataset Language Signers Videos Duration (hours) Vocab Resolution FPS Source
SIGNUM[14] German 25 11,375 8.43 455 $776\times 578$ 30 Lab
PHOENIX-2014[11] German 9 6,841 10.71 1,231 $210\times 260$ 25 TV
PHOENIX-2014T[12] German 9 8,257 10.53 1,231 $210\times 260$ 25 TV
GSL[13] Greek 7 10,295 9.59 310 $848\times 480$ 30 Lab
CSL[15] Chinese 50 25,000 100+ 178 $1280\times 720$ 30 Lab
CSL-Daily[16] Chinese 10 20,654 23.27 2,000 $1920\times 1080$ 30 Lab
TVB-HKSL-News[17] Chinese 2 7,160 16.07 6,515 $248\times 360$ 25 TV
CE-CSL(ours) Chinese 12 5,988 10.52 3,515 varying varying Lives
@@ II-A Continuous sign language dataset
A CSL dataset serves as an indispensable cornerstone for researching CSLR technology, with its quality directly determining the performance and application potential of recognition algorithms[16][18]. In recent years, with the rapid development of artificial intelligence technologies, a series of CSL datasets have emerged, contributing significantly to the advancement of this field. Agris et al.[14] pioneered the creation of the SIGNUM sign language dataset, which includes rich samples of isolated words and continuous sentences performed by 25 signers of different genders and age groups. Although SIGNUM made initial explorations in diversity and corpus size, the limitation of 450 different German sign language words and 780 continuous sentences did not adequately meet the large-scale data requirements of deep learning algorithms. In 2015, Koller et al.[11] introduced the PHOENIX14-2014 dataset, which collected 6,841 continuous sentences through weather forecast programs on German public television, significantly enhancing the scale and practicality of the dataset. Building on this, Camgoz et al.[12] extended the dataset to create the PHOENIX14-2014T dataset, enhancing its professionalism and introducing refined sign boundary definitions, annotated by deaf experts, further improving the dataset’s quality. Huang et al.[15] focused on Chinese sign language, collecting video data including RGB, depth, and body joint patterns using Microsoft Kinect cameras, providing a multimodal perspective for sign language recognition research. Adaloglou et al.[13] concentrated on practical application scenarios of sign language, building cases involving interactions with police departments, hospitals, and citizen service centers, enriching the diversity of scenes in the dataset. Niu et al.[17] utilized an automated data collection pipeline to systematically gather extensive continuous sign language videos from seven months of sign language news broadcasts by Hong Kong Television Broadcasts Limited (TVB), further expanding the publicly available data resources. Despite the critical role these datasets have played in advancing sign language recognition technology, most originate from laboratory environments or television programs, commonly featuring a single background and uniform lighting conditions, which differ significantly from real-life sign language communication scenarios. In Table I, we list some large-scale publicly available continuous sign language datasets based on video. Given the limitations of existing datasets, we have constructed a Chinese CSL dataset that truly reflects the complexity of real-life environments. Unlike previous datasets limited to laboratory settings or television programs, ours involves real-world collection by sign language performers, aiming to capture the dynamics of sign language in authentic scenarios, including varied backgrounds, natural lighting conditions, and rich contextual changes.
II-B Continuous sign language recognition
In the field of CSLR, precise frame-level annotation has always been an insurmountable challenge, with its complexity and time-consuming nature limiting the scale and quality of datasets[19]. As a result, most research has focused on video-level annotation, treating CSLR as a weakly supervised learning problem, where there is no strict one-to-one correspondence between video frames and annotation sequences, adding to the difficulty of the recognition task[20][21][22]. However, with the rise of deep learning technologies, CSLR has seen revolutionary advancements[23][24][25]. The introduction of Convolutional Neural Network (CNN) has significantly improved the extraction of frame-level sign language features, while Recurrent Neural Network (RNN) have complemented static feature representations with their superior ability to handle temporal sequences, together forming the basic framework of current CSLR systems.
Hu et al.[26] innovatively introduced a correlation module capable of dynamically computing the correlation mapping between the current frame and adjacent frames, effectively identifying the trajectories of spatial patches, providing a new perspective for CSLR. Shi et al.[27] proposed an efficient training strategy that successfully transferred high-quality visual features to the task of continuous sign language recognition, significantly enhancing the recognition accuracy of the model. Zhu et al.[28] focused on dynamic attention mechanisms, which can effectively capture the nonlinear changes in local motion regions during the expression of sign language, generating dynamic descriptions of image changes, thereby further strengthening the model’s ability to capture complex sign language dynamics. Gao et al.[29], drawing inspiration from enhancement schemes in signal processing, innovatively designed a Temporal Lifting Pool, decomposing input signals into multiple sub-bands, each corresponding to a specific motion pattern, effectively enhancing the extraction capabilities of temporal features. Li et al.[30] dedicated themselves to constructing a multi-scale temporal feature network, which, through time receptive fields at different scales, significantly improved the flexibility and accuracy of sequence modeling by the model.
Addressing the challenges posed by complex backgrounds to sign language recognition, we propose a TFNet model for CSLR. This method aims to achieve efficient and accurate sign language recognition by extracting sequence features in both the temporal and frequency domains, opening up new avenues for the development of CSLR technology.
@@ II-B Continuous sign language recognition
TABLE II: Statistical Data of the CE-CSL Dataset.(OOV: out-of-vocabulary, e.g., words that occur in Dev set or Dev set but not in Train set.)
@@
A 490 55 55 iPhone 14 Plus $20-25$ $\operatorname{man}$
B 490 53 57 iPhone 12 Pro $20-25$ woman
C 504 54 42 iPhone 13 $20-25$ man
D 507 46 47 iPhone 13 $20-25$ woman
E 495 51 54 iPhone 13 $20-25$ $\operatorname{man}$
F 492 50 58 HUAWEI P50 Pro $25-30$ woman
G 490 54 56 OnePlus Ace $25-30$ woman
H 101 9 9 iPhone 13 Pro $25-30$ $\operatorname{man}$
I 66 9 9 iPhone 13 Pro $40-45$ woman
J 330 29 38 iPhone 14 Pro Max $20-25$ woman
K 501 53 46 iPhone 12 $20-25$ woman
L 507 52 29 iPhone 14 Pro $20-25$ woman
Figure 2: Illustration of Backgrounds in Selected Images from the CE-CSL Dataset.
@@ III Dataset and Methodology
III-A CE-CSL dataset
To promote the transition of CSLR technology from laboratory environments to real-life applications, we have carefully constructed a publicly available Chinese Continuous Sign Language dataset named CE-CSL. Unlike the limitations of previous datasets, CE-CSL aims to simulate natural living scenarios, without restricting recording equipment, attire choices, or background environments, ensuring that the collected sign language videos closely resemble real-world application contexts.
The CE-CSL dataset encompasses 12 sign language performers (labeled A through L), including eight females and four males. Notably, performers I and J are hearing-impaired individuals, while the remaining performers are professional sign language interpreters, ensuring a diverse and authentic representation of sign language expressions within the dataset. The dataset comprises a total of 5,988 continuous sign language videos. According to strict division ratios, there are 4,973 sign language videos in the training set, 515 in the validation set, and 500 in the test set. These videos collectively involve 3,515 Chinese words, covering a broad range of daily communication needs. The video lengths vary widely, ranging from the shortest at 39 frames to the longest at 530 frames, showcasing the richness and flexibility of sign language expression. Detailed statistics of the CE-CSL dataset are presented in Table II.
We also provide a comprehensive listing of key parameters for each performer in Table III, including the recording equipment used, the number of videos, age, and gender, offering detailed baseline information for future research. Given the open setting for background environments, the CE-CSL dataset includes over 70 different living scenarios, greatly enriching the diversity of the data. To better understand the diversity of backgrounds in the CE-CSL dataset, we present a series of representative scenes in Figure 2, including indoor, outdoor, park, street, mall, office, and others, all sourced from our dataset. Through these images, one can clearly see the extensive range of scenarios covered by the CE-CSL dataset, further demonstrating its valuable contribution as a research resource and its significance in advancing the practical application of continuous sign language recognition technology.

### v2

@@ Abstract
The current bottleneck in continuous sign language recognition (CSLR) research lies in the fact that most publicly available datasets are limited to laboratory environments or television program recordings, resulting in a single background environment with uniform lighting, which significantly deviates from the diversity and complexity found in real-life scenarios. To address this challenge, we have constructed a new, large-scale dataset for Chinese continuous sign language (CSL) based on complex environments, termed the complex environment - Chinese sign language dataset (CE-CNSL). This dataset encompasses 5,988 continuous CSL video clips collected from daily life scenes, featuring more than 70 different complex backgrounds to ensure representativeness and generalization capability. We also propose a time-frequency sequence based continuous sign language recognition model (TFNet) to address the issue of advanced CSLR methods ignoring frequency domain features. After frame level feature extraction, this model utilizes both time-domain and frequency-domain information to extract sequence features separately, and then fuses them to achieve efficient and accurate continuous sign language recognition. Experimental results demonstrate that our approach achieves significant performance improvements on the CE-CNSL, validating its effectiveness under complex background conditions. Additionally, our proposed method has also yielded highly competitive results when applied to three publicly available CSL datasets. Code and dataset is available at https://github.com/woshisad159/TFNet.git.
@@ I Introduction
Sign language recognition has always been a key issue of concern for researchers. As a model of interdisciplinary research, it integrates the essence of computer vision, natural language processing, and human-computer interaction technology[1][2]. Sign language recognition technology not only simplifies the complexity of cross linguistic communication by translating sign language dynamic videos into understandable text or speech in real time, but also demonstrates enormous potential in multiple fields such as education, employment, and public services[3].
@@
In the field of CSLR, achieving accurate recognition of continuous sign language relies heavily on having a high-quality CSL dataset[4][5][6][7]. There are currently four commonly used publicly available sign language recognition datasets. Koller et al.[8] constructed the PHOENIX14-2014 dataset by collecting weather forecast programs broadcasted by a German public television station, providing valuable resources for CSLR research. This dataset is based on demonstrations by professional sign language interpreters; however, its collection environment remains limited to the television studio, lacking the diversity of real-world contexts. Camgoz et al.[9] further expanded upon this by introducing the PHOENIX14-2014T dataset, increasing the volume of data and refining the definition of sign boundaries, thereby enhancing annotation accuracy. Although this dataset improves in terms of professionalism, it is still constrained by relatively idealized background conditions, making it difficult to comprehensively reflect the sign language communication scenarios of everyday life. Adaloglou et al.[10], focusing on practical application scenarios, particularly emphasized communication between deaf individuals and public service institutions, such as police departments, hospitals, and citizen service centers. They meticulously designed five typical scenarios, simulated dialogues of varying difficulty levels, thus enhancing the practicality and authenticity of the dataset. Zhou et al.[11] constructed a large-scale, continuous dataset of Chinese sign language, focusing on people’s daily life scenarios. This dataset is based on demonstrations by professional sign language interpreters and covers multiple topics such as family life, healthcare, and school life. Although the above-mentioned datasets have played an important role in advancing continuous sign language recognition technology, they generally face a common challenge: the collection environment is disconnected from real life, making it difficult to fully reflect sign language communication scenarios in daily life, as shown in Figure 1.
@@
Figure 1: The above are partial image sequences from four publicly available continuous sign language datasets, PHOENIX14-2014, PHOENIX14-2014T, GSL, and CSL-Daily. It can be seen from these image sequences that these datasets were collected in laboratory environments or television programs, with the background of the subjects being a solid color background and uniform lighting
@@
To overcome the limitations of existing continuous sign language datasets, particularly their disconnection from real-life scenarios, this study constructs a CSL dataset oriented towards practical application environments. The dataset aims to facilitate the seamless transition of CSLR technology from laboratory settings to everyday life, laying a solid foundation for barrier-free communication between the deaf and hearing communities. The construction process of the CE-CNSL dataset strictly adheres to the principle of practical application orientation, collecting a large number of continuous sign language video materials derived from real-life scenarios, covering a wide range of situational variations and environmental complexities. After the completion of the CE-CNSL dataset, we conducted a comprehensive evaluation using advanced deep learning models to verify its reliability and value as a research benchmark.
Figure 2: Heat maps of feature vector sequences in both time and frequency domains for a sign language video after frame level feature extraction
In addition, temporal feature extraction is a crucial step in CSLR tasks. Sign language, as a highly dynamic visual language, relies not only on the static shape of gestures, but also on the motion trajectory, temporal information, and dynamic features of gestures. These dynamic features change over time and can describe the evolution of sign language expression. However, the current method [12][13][14] only considers the temporal aspect and ignores the characteristics in the frequency domain. The frequency domain based sign language features can provide another perspective on the periodic motion patterns and rhythm changes in sign language. Different feature forms can greatly enrich the expression of sign language features, as shown in Figure 2.
This paper proposes a new sequence feature extraction module based on the consideration of feature diversity. By using 1DCNN BILSTM for sequence feature modeling in both time and frequency domains, rich sign language feature expressions are obtained by integrating information from different domains. An end-to-end continuous sign language recognition model based on time-frequency sequences (TFNet) is also proposed. This combination of cross domain features not only enriches the model’s feature information, but also enhances its ability to recognize complex patterns, thereby demonstrating higher performance in various application scenarios. The experimental results demonstrate that TFNet not only achieves highly competitive results on currently publicly available datasets, but also shows significant performance improvements compared to existing advanced models on the CE-CNSL dataset.
@@
    We have constructed a Chinese Continuous Sign Language dataset CE-CNSL dataset oriented towards practical application environments. This dataset, with its rich content and precise annotations, aims to provide a robust foundational resource for the advancement of CSLR technology.
@@
    We propose a new sequence feature extraction module that innovatively extracts frequency domain features. By integrating different domain feature information, enriching feature expression, and obtaining efficient and accurate semantic expression, it provides a new perspective for exploring the periodic motion patterns and rhythm changes in sign language. It can effectively reveal the imperceptible periodic patterns in sign language actions and significantly improve the accuracy of continuous sign language recognition in complex environments.
@@
    We propose a continuous sign language recognition model based on time-frequency sequences (TFNet), which greatly expands the dimensions of comprehensive understanding of sign language features, providing indispensable support for a deeper understanding of the prosodic structure of sign language and improving the accuracy of sign language recognition. It has high universality and robustness on different continuous sign language recognition datasets.
@@
Dataset Language Signers Videos Duration (hours) Vocab Resolution Modalities FPS Source
SIGNUM[15] German 25 11,375 8.43 455 $776\times 578$ RGB 30 Lab
PHOENIX-2014[8] German 9 6,841 10.71 1,231 $210\times 260$ RGB 25 TV
PHOENIX-2014T[9] German 9 8,257 10.53 1,231 $210\times 260$ RGB 25 TV
GSL[10] Greek 7 10,295 9.59 310 $848\times 480$ RGB 30 Lab
CSL[16] Chinese 50 25,000 100+ 178 $1280\times 720$ RGB+D 30 Lab
CSL-Daily[11] Chinese 10 20,654 23.27 2,000 $1920\times 1080$ RGB+D 30 Lab
TVB-HKSL-News[17] Chinese 2 7,160 16.07 6,515 $248\times 360$ RGB 25 TV
CE-CNSL(ours) Chinese 12 5,988 10.52 3,515 varying RGB varying Lives
@@ II-A Continuous sign language dataset
A CSL dataset serves as an indispensable cornerstone for researching CSLR technology, with its quality directly determining the performance and application potential of recognition algorithms[11][18]. In recent years, with the rapid development of artificial intelligence technologies, a series of CSL datasets have emerged, contributing significantly to the advancement of this field. Agris et al.[15] pioneered the creation of the SIGNUM sign language dataset, which includes rich samples of isolated words and continuous sentences performed by 25 signers of different genders and age groups. Although SIGNUM made initial explorations in diversity and corpus size, the limitation of 450 different German sign language words and 780 continuous sentences did not adequately meet the large-scale data requirements of deep learning algorithms. In 2015, Koller et al.[8] introduced the PHOENIX14-2014 dataset, which collected 6,841 continuous sentences through weather forecast programs on German public television, significantly enhancing the scale and practicality of the dataset. Building on this, Camgoz et al.[9] extended the dataset to create the PHOENIX14-2014T dataset, enhancing its professionalism and introducing refined sign boundary definitions, annotated by deaf experts, further improving the dataset’s quality. Huang et al.[16] focused on Chinese sign language, collecting video data including RGB, depth, and body joint patterns using Microsoft Kinect cameras, providing a multimodal perspective for sign language recognition research. Adaloglou et al.[10] concentrated on practical application scenarios of sign language, building cases involving interactions with police departments, hospitals, and citizen service centers, enriching the diversity of scenes in the dataset. Niu et al.[17] utilized an automated data collection pipeline to systematically gather extensive continuous sign language videos from seven months of sign language news broadcasts by Hong Kong Television Broadcasts Limited (TVB), further expanding the publicly available data resources. Despite the critical role these datasets have played in advancing sign language recognition technology, most originate from laboratory environments or television programs, commonly featuring a single background and uniform lighting conditions, which differ significantly from real-life sign language communication scenarios. In Table I, we list some large-scale publicly available continuous sign language datasets based on video.
Given the limitations of existing datasets, we have constructed a Chinese continuous sign language dataset that truly reflects the complexity of real life. Unlike being limited to laboratories or television programs in the past, our dataset is collected by sign language presenters in daily environments, aiming to capture sign language dynamics in real-life scenarios, including changing backgrounds, natural lighting conditions, and rich contextual changes.
@@
TABLE II: Statistical Data of the CE-CNSL Dataset.(OOV: out-of-vocabulary, e.g., words that occur in Dev set or Test set but not in Train set.)
@@
A 490 55 55 iPhone 14 Plus 202520-25 $\operatorname{man}$
B 490 53 57 iPhone 12 Pro 202520-25 woman
C 504 54 42 iPhone 13 202520-25 man
D 507 46 47 iPhone 13 202520-25 woman
E 495 51 54 iPhone 13 202520-25 $\operatorname{man}$
F 492 50 58 HUAWEI P50 Pro 253025-30 woman
G 490 54 56 OnePlus Ace 253025-30 woman
H 101 9 9 iPhone 13 Pro 253025-30 $\operatorname{man}$
I 66 9 9 iPhone 13 Pro 404540-45 woman
J 330 29 38 iPhone 14 Pro Max 202520-25 woman
K 501 53 46 iPhone 12 202520-25 woman
L 507 52 29 iPhone 14 Pro 202520-25 woman
II-B Continuous sign language recognition
Effectively extracting and utilizing temporal features is of great significance for improving the accuracy and robustness of continuous sign language recognition[19][20][21]. Hu et al.[22] innovatively introduced a correlation module that can dynamically calculate the correlation mapping between the current frame and adjacent frames, effectively identifying the trajectory of spatial patches and providing a new perspective for continuous sign language recognition. Zhu et al.[23] focused on the dynamic attention mechanism, which can effectively capture the nonlinear changes in local motion areas during sign language expression, generate dynamic descriptions of image changes, and further enhance the model’s ability to capture complex sign language dynamics. The above methods mainly focus on the motion trajectory between images, and improve the accuracy of sign language recognition by enhancing the ability to capture dynamic features. In terms of temporal feature extraction, only a hybrid convolutional neural network composed of 1DCNN and BiLSTM is used for feature extraction. Li et al.[24] are committed to building a multi-scale temporal feature network, which significantly improves the flexibility and accuracy of the model in modeling sequences through different scales of temporal receptive fields. However, the introduction of multi-scale convolution kernels resulted in a sharp increase in the number of model parameters. Zhou et al.[25] proposed a Spatiotemporal Multi Clue (STMC) network to solve vision based sequence learning problems. The network consists of a Spatial Multi Clue (SMC) module and a Temporal Multi Clue (TMC) module. The introduction of multiple clues greatly improves the accuracy of the model, but it is also the use of multiple clues that increases the computational complexity and inference time of the model. These methods, like many existing methods[26][27][28][29], globally pool the extracted frame level features, convert them into one-dimensional sequence features related to time, and then perform sequence feature extraction. Only considering the temporal aspect, while ignoring the characteristics in the frequency domain. This paper proposes a continuous sign language recognition model based on time-frequency sequences (TFNet). This method aims to achieve efficient and accurate sign language recognition by jointly extracting sequence features in both time and frequency domains, opening up a new path for the development of continuous sign language recognition technology.
Figure 3: Illustration of backgrounds in selected images from the CE-CNSL dataset.
@@ III Dataset and Methodology
III-A CE-CNSL dataset
To promote the transition of CSLR technology from laboratory environments to real-life applications, we have carefully constructed a publicly available Chinese Continuous Sign Language dataset named CE-CNSL. Unlike the limitations of previous datasets, CE-CNSL aims to simulate natural living scenarios, without restricting recording equipment, attire choices, or background environments, ensuring that the collected sign language videos closely resemble real-world application contexts.